Learning Tasic No. 1


A die is lossed. What is the probability of getting


1.3 _____


2. An even number _____


3.7 _____


4. An odd number _____


5. A prime number _____


1
Expert's answer
2022-05-30T17:00:17-0400

1.


"P(3)=\dfrac{1}{6}"

2.


"P(even)=\dfrac{3}{6}=\dfrac{1}{2}"

3.


"P(7)=\dfrac{1}{6}"

4.


"P(odd)=\dfrac{3}{6}=\dfrac{1}{2}"

5.

Prime numbers: "1,2,3,5"


"P(prime)=\dfrac{4}{6}=\dfrac{2}{3}"